40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Wed May 1 15:40:46 2019, Koji, Optics, Characterization, OMC(004): Spot positions and the scattering misalignment.pdf
    Reply  Fri May 3 11:06:28 2019, Koji, Optics, Characterization, OMC(004): Spot positions and the scattering misalignment.pdf
Message ID: 357     Entry time: Fri May 3 11:06:28 2019     In reply to: 356
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: OMC(004): Spot positions and the scattering 

Experiment on 5/1
- CM1 right knob was moved 1div (10um) backward such that the spots were better centered on the mirrors 

FM1 (A5): h=-0.2mm -> 0.4mm made the spot much darker but still it has a few scattering spots.
FM2 (A14): h=-0.8mm -> 0.2mm reduced the number of spots from 2 to 1. And it is darker. The remaining spot at the center.
CM1 (C11): h=-1.3mm -> +1.0mm made the spot much darker.
CM2 (C12): h=-0.7mm -> +0.2mm remains dark.

Note: CM1 h=1mm and CM2 h~0mm are good locations. h+ is the good direction to move. Avoid h-.
FM1 and FM2 has the scat spots at the center. Want to go h+ more.

Uniformly go h+ is the good move. => This can be done by rotate CM1 positive => CM1 right knob CCW.

2019/5/1 CM1 right micrometer 1div backward
         
    Unit   V_RefPD [V]
P_TRANS 13.53 [mW]   3.09
V_REFL_LOCKED 53.4 [mV]   3.09
V_REFL_UNLOCK 2.52 [V]   3.065
P_IN 14.45 [mW]   3.07
V_REFL_OFFSET -6.35 [mV]    
         
Coupling 0.977      
OMC_Trans 0.953      

Improvement of the transmission from 93.9%->95.3%


- Further moved CM1 right knob 0.5div (0.5um) backward such that the spots were moved to h+ directions.
FM1 (A5): h=0.4mm -> 1.1mm (there is only one spot rather than multiple spots)
FM2 (A14): h=0.2mm -> 1.1mm (darker but multiple spots)
CM1 (C11): h=1.0mm -> 1.8mm (brighter but single spot)
CM2 (C12): h=0.2mm -> 1.5mm (dark multiple spots)

2019/5/1 CM1 right micrometer 0.5div backward
         
    Unit   V_RefPD [V]
P_TRANS 14.55 [mW]   3.28
V_REFL_LOCKED 49 [mV]   3.28
V_REFL_UNLOCK 2.755 [V]   3.299
P_IN 15.64 [mW]   3.3
V_REFL_OFFSET -6.316 [mV]    
         
Coupling 0.980      
OMC_Trans 0.955      

Not much improvement of the transmission but kept 95% level.

- Replaced FM1 (A5) with A1 mirror (No photo)

Good news: This did not change the cavity alignment at all.

Transmission 95.4%

- Tweaked the CM1 angle

Transmission 95.3%

=> A1 mirror does not improve the transmission much.


Next Plan: Use A5 (or something else) as FM2 and see if A14 caused the dominant loss.

Attachment 1: misalignment.pdf  18.026 MB  Uploaded Fri May 3 13:00:44 2019  | Hide | Hide all
misalignment.pdf misalignment.pdf misalignment.pdf
ELOG V3.1.3-